Solve (2D^2D'-3DD'^2+D'^3)z=0
1
Expert's answer
2020-11-10T19:43:44-0500
"Solution"

Given "(2D^2D'-3DD'^2+D'^3)z=0"


"D'(D^2-\frac32DD'+\frac12D'^2)z=0\\
D'(D^2-\frac12DD'-DD'+\frac12D'^2)z=0 \implies D'(D(D-\frac12D')-D'(D-\frac12D))\\
\therefore (D'(D-D')(D-\frac12D'))z=0"


Then

"D'\implies\phi_1(y)\\
(D-D')\implies\phi_2(y+x)\\
(D-\frac12D')\implies\phi_3(y+\frac12x)\\

z= \phi_1(y)+\phi_2(y+x)+\phi_3(y+\frac12x)\\"

Answer


"z= \phi_1(y)+\phi_2(y+x)+\phi_3(y+\frac12x)\\"